OpenStudy (anonymous):

Question is posted as a comment below:

OpenStudy (saifoo.khan):

@owl, did u saw the chat?

OpenStudy (anonymous):

\[Solve x/21\le-6\]

OpenStudy (saifoo.khan):

\[\frac{x}{21} \le -6\]Multiply 21 on both sides, \[x \le -6 \times 21\] \[x \le -126\]

OpenStudy (anonymous):

Ok. Thanks! :)

OpenStudy (anonymous):

Help me with this one: Solve: \[-x/4\le\]

OpenStudy (anonymous):

Oh there's a 7 after the sign

OpenStudy (lilg132):

so in this case all you do is whatever x is being divided by to get rid of it you must multiply both sides by that number so \[x / 21 \le - 6\] now you multiply both sides by what x is being divided by to get x by it self and also to solve x so \[x / 21 * 21 \le - 6 * 21\] which will give you \[x \le -126\]

OpenStudy (anonymous):

I don't think that's right because that's not one of the choices on my homework.

OpenStudy (anonymous):

I'll post the equation again so it's less confusing.

OpenStudy (anonymous):

\[-x/4\le\]

OpenStudy (lilg132):

this is another inequality question? by the way that is incomplete?

OpenStudy (anonymous):

Hmm. I wonder why it's not putting the 7.

OpenStudy (lilg132):

oh you have to wait about 2 or 3 seconds after you put 7

OpenStudy (lilg132):

then click insert

OpenStudy (anonymous):

oh ok. I'll try it.

OpenStudy (anonymous):

\[-x/4\le7\]

OpenStudy (anonymous):

There we go.

OpenStudy (lilg132):

this is a different question right?

OpenStudy (anonymous):

I think so...

OpenStudy (lilg132):

ok did you have a read through the method i posted above its the same steps but in the end you should have an answer you may be stuck with when you get that i will help you further

OpenStudy (lilg132):

so try and follow my steps and see if you can manage to do that if not let me know where you are stuck